If $2000 is invested at 8% interest, how much must be invested at 11% interest so that the total return for both averages are 10%?
Answer by mananth(16946)

2000------------------------8%
x---------------------------11%
(x+2000)-------------------10%

2000*8%+11%x= (x+2000)*10%

multiply by 100
8*2000+11x+10(x+2000)
16000+11x=10x+20000
11x-10x= 20000-16000
x=4000